OpenStudy (anonymous):

A particle is moving along a projectile path where the initial height is 160 feet with an initial speed of 144 feet per second. What is the maximum height of the particle? Answer 224 feet 272 feet 128 feet 484 feet

OpenStudy (anonymous):

Please post only relevant question in that group. I am closing this now. -Thanks.

OpenStudy (anonymous):

this is math

OpenStudy (anonymous):

using quadratic functions for real life scenarios

OpenStudy (anonymous):

Oh Crap! Mea culpa. Okay as I have closed this one, I will help you with this one.

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

I can see why but were not using a physics forumla I believe... Quadratic functions are math idk about physics

OpenStudy (anonymous):

Okay, what have you tried so far?

OpenStudy (anonymous):

Assuming Earth's gravity, the formula for the flight of the particle is: s(t) = -16t^2 + vt + s = -16t^2 + 144t + 160. This has a maximum when t = -b/(2a) = -144/(-32) = 9/2. Therefore, the maximum height is s(9/2) = -16(9/2)^2 + 144(9/2) + 160 = 484 ft. I hope this helps!

OpenStudy (anonymous):

I was told if I used this forumla and plug in the variables it would work s(t) = -16t^2 + 144t + 160

OpenStudy (anonymous):

o wow then he posts it while I was typing it

OpenStudy (anonymous):

I'm just bad at plugging in variables is all

OpenStudy (anonymous):

ZhangYan literally solved it for you ;)

OpenStudy (anonymous):

thanks ffm